I am having a rough time solving my problems on probability and standard deviation. One problem in particular, Mean is 12.29 and standard deviation is .09. The problem asks for the percentage of cans that contain less than 12 ounces. I'm really not sure even where to start with this one. Please help.
Shanen

Answer by stanbon(57352)

Mean is 12.29 and standard deviation is .09. The problem asks for the percentage of cans that contain less than 12 ounces.
---------------
Convert 12 to a z-score:
z(12) = (12-12.29)/0.09 = -0.29/0.09 = -3.222222...
---------
Now, P(x<12)=P(z<-3.22222...) = 0.000636...
In percentage this is 0.06%
===================
Cheers,
Stan H.